# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

### v1

@@ 4.3 Results and Discussion
Figure 2 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as a function of $N$ for three text strings generated by the same GPT-2 model using top-k sampling with $k=100$ for Fig. 2(a) and 2(b) and $k=4$ for Fig. 2(c). The plots also shows a range around the empirical entropy corresponding to ±$\pm\lambda_{N}({\mathbf{Y}}_{N})$ and ±$\pm 2\lambda_{N}({\mathbf{Y}}_{N})$. We see that in all three cases, the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently. We also see that the long-term average empirical entropy is around $3.5$ bits per token in Fig. 2(a) and Fig. 2(b), and $1.5$ bits per token in Fig. 2(c), which corresponds to approximately $2^{3.5}\approx 11$ alternative tokens to choose from for top-100 sampling and $2^{1.5}\approx 3$ tokens for top-4 sampling.
(a) GPT-2-generated string with top-100 sampling.
(b) GPT-2-generated string with top-100 sampling.
(c) GPT-2-generated string with top-4 sampling.
Figure 2: Log-perplexity and empirical entropy for GPT-2 of strings $\mathbf{Y}$ generated by the same model.
(a) Human written text.
(b) Claude 3 written text.
(c) Decreasing entropy.
@@
Figure 3: Log-perplexity and empirical entropy for GPT-2 of selected text strings.
Figures 3(a) and 3(b) show the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as described in Section 4.2 for a string not generated by the GPT-2 model. The string in Fig. 3(a) is a paragraph written by the authors for an early draft of this manuscript, and the string in Fig. 3(b) was generated by the Claude 3 Opus model from Anthropic (2024). While our theory does not require non-convergence of $l_{M}({\mathbf{Y}}_{N})$ to $h_{M}({\mathbf{Y}}_{N})$, we do see from Fig. 3(a) and Fig. 3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ away from the empirical entropy $h_{M}({\mathbf{Y}}_{N})$. The contrast between Figs. 3 and 2 has obvious implications for AI text detection (Tang et al., 2024), model fingerprinting (Yu et al., 2021) and so on.
4.4 Anomalies and Limitations
While the results from our experiments described in Section 4 were always consistent with Proposition 3, they did reveal certain limitations of our setup that we now discuss.
For all three strings in Fig. 2, we see that $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$, but the log-deviations $\lambda_{M}({\mathbf{Y}}_{N})$ fail to become negligibly small even at $N=512$ suggesting that the GPT-2 model is not capable of generating strings sufficiently long for asymptotics to apply.
Fig. 3(c) shows an example of an output generated by GPT-2 for which the log-perplexity and empirical entropy both monotonically decrease with string length over many tokens. This suggests that the model repeatedly produces very low entropy token distributions $p_{n}(y)$, which happens for degenerate distributions where the probability mass is concentrated in a small number of tokens. In fact, on closer examination, we find that in these cases, the model is stuck in a state where the next token distribution just picks one token deterministically. This results in output text that gets stuck in a loop, repeating the same substring until $N_{max}$ has been reached.
Fundamentally, these limitations arise from cost and computational constraints. More powerful LLMs can generate longer texts and be less likely to get stuck in bad states that produce degenerate distributions and so on. In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti). Additional plots and more details about these experiments are presented in Section A.1.
@@ 5 Conclusions
We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property. While this paper focused narrowly on an exposition of the AEP property, it opens up many questions for further study. For instance, Proposition 4 suggests that different language models have distinct and unique “model signatures.” The contrast between Fig. 2 and Fig. 3 suggests potential applications for understanding and enhancing AI detection algorithms. As reported in Section 4.4, our experimental results are limited by computational constraints and can be strengthened by improved hardware and software. A broader question is to explore other statistical laws and constraints that generative models are bound by.
Appendix A Supplemental Material
A.1 Additional Experimental Context and Results
In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti). On this setup, each run of a text generation (i.e., running GPT-2 to generate $N=512$ tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds. We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.
Figure 4 shows additional results similar to Fig. 2 where we naively varied the seed value of the random number generator from 100-105. We present these to show that Fig. 2’s results are representative of the general case and were not cherry-picked. In all cases, $K=100$ and $N=512$. Further, should the community like to validate or expand upon this work, we have open sourced our repository [Bell and Mudumbai, 2024]. In addition to the source code, the supplemental material includes an /examples directory that includes the target text and the anomalous generated text for the plots in Fig. 3.
(a) Seed=100.
(b) Seed=101.
(c) Seed=102.
(d) Seed=103.
(e) Seed=104.
(f) Seed=105.
Figure 4: Log-perplexity and empirical entropy for GPT-2-generated string with top-100 sampling.
A.2 A Simple Markov Language Model

### v3

@@ 4.3 Results and Discussion
Figure 2 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 or Llama 3.1 8B model as a function of $N$ for several text strings generated by the same model. The plots also shows the range $h_{M}({\mathbf{Y}}_{N})\pm\lambda_{M}({\mathbf{Y}}_{N})$ and $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$. We see that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently in all cases. The asymptotic convergence is more clearly seen in the long strings from the Llama 3.1 8B model [^5] The GPT-2 model cannot generate similarly long texts. in Fig. 2(c), 2(d) . We find that very long synthetic texts such as those in Fig. 2(c), 2(d) can be ungrammatical, repetitive and generally of poor quality, as reported elsewhere (Holtzman et al., 2020). However, Proposition 3.1 holds regardless of the quality of the text as seen in Fig. 2.
(a) GPT-2, $N_{max}=512$.
(b) Llama 3.1 8B, $N_{max}=1024$.
(c) Llama 3.1 8B, $N_{max}=10240$.
(d) Llama 3.1 8B, $N_{max}=10240$.
@@
Figure 2: Log-perplexity and empirical entropy for GPT-2 or Llama 3.1 8B of self-generated text, top-100 sampling.
Figure 3 shows the log-perplexity and empirical entropy for the Llama 3.1 8B model for two long, non-Llama generated strings. The text used in Fig. 3(a) is an excerpt from a recent article (Pelly, 2025) in Harper’s Magazine, a respected literary publication. The string used in Fig. 3(b) was generated by ChatGPT o1 Pro (OpenAI, 2024).
(a) Human writing.
(b) Text from ChatGPT o1 pro.
Figure 3: Log-perplexity and empirical entropy for Llama 3.1 8B of non-Llama-generated strings $\mathbf{Y}$.
We see from Fig. 3(a) and Fig. 3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ for Llama 3.1 8B is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ larger than the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ showing that the corresponding strings are clearly under-typical for the Llama 3.1 8B model. While this non-convergence is not required by our theory - we have not considered whether the converse to Proposition 3.1 is true - it is strongly suggested by Propositions 3.3 and 3.4 if we assume that the strings in Fig. 3 are randomly sampled from the set of all grammatically correct texts.
4.4 More Catholic than the Pope
(a) Alice in Wonderland excerpt.
(b) The Great Gatsby excerpt.
Figure 4: Log-perplexity and empirical entropy for Llama 3.1 8B of excerpts from literary classics.
This brings us to the remarkable plots in Fig. 4 which shows the log-perplexity of excerpts from classic texts for Llama 3.1 8B, specifically from Alice in Wonderland and The Great Gatsby in Figs. 4(a) and 4(b) respectively. The average entropy for the token distributions in Fig. 4(a) is an astonishingly small $0.11$ bits, which means that the model on average assigns a very high probability ($\sim 99\%$) to the precise sequence of tokens from the excerpt for its predicted next tokens. In other words, this excerpt has been memorized by the Llama 3.1 8B model. The same observation also applies to Fig. 4(b). We also see from both plots that the log-perplexity is several standard deviations below the empirical entropy making the excerpts slightly, but unambiguously, over-typical for the Llama-8B model: the model judges these texts to be more perfect than its own creations!
From Proposition 3.4, we can conclude that these excerpts are among the small number of very special strings to be so memorized by the model. Clearly, these texts were part of the training set of the Llama-8B model and they were esteemed highly by its cost function.
Interestingly, the over-typicality of these excerpts means that it is pretty much impossible for the Llama-8B model itself to generate them verbatim - even with a $99\%$ probability that each next token will exactly match the excerpt, a randomly generated string can be expected to diverge from the excerpt after only about $100$ tokens.
Additional plots and more details about these experiments are presented in Appendix C.1.
5 Practical Applications
Consider the problem of AI text detection i.e. determining whether a string ${\mathbf{Y}}$ was generated by a model M. The intuition behind (9) suggests a simple method for doing this: check if the perplexity is small enough. Specifically, declare ‘yes’ if $l_{M}({\mathbf{Y}})<l_{th}$ for a threshold $l_{th}$. The same test can also be used for the problem of LLM dataset inference i.e. determining whether a string ${\mathbf{Y}}$ was used as part of the training dataset for the model M.
While, this idea is not new and indeed variations of it are behind popular solutions for both problems (Mitchell et al., 2023; Maini et al., 2024), our results in this paper provide a rigorous theoretical foundation for it. It also provides a principled way to choose the classification threshold $l_{th}$ which previously had to be computed empirically (Carlini et al., 2021). As an example, with $l_{th}=h_{M}({\mathbf{Y}})+3\lambda_{M}({\mathbf{Y}})$, Fig. 2 shows that we can accurately distinguish between synthetic text generated by Llama 3.1 8B and non-Llama text such as in Fig. 3 even with relatively short strings of $1000$ tokens. We can also provide performance guarantees. Indeed, a simple lower-bound on the probability of a false negative (i.e. failing to detect a Llama-generated string) can be obtained from (18). Similar comments apply to the dataset inference problem: while it is well-known that text strings memorized by an LLM have low perplexity, our theory allows us to rigorously derive threshold values and performance guarantees. Extrapolating from Fig. 4, we posit that if a long excerpt from a human-written book is over-typical for a language model, it is a near-statistical certainty that the book was part of the training data set for that model.
@@ 6 Conclusions
We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property. While this paper focused on a rigorous exposition of the AEP property, it opens up many questions for further study. An important open problem is to design practical algorithms for AI text detection and dataset inference using the typical set concept. Tighter non-asymptotic bounds on deviations that improve on (18) will significantly benefit practical applications by providing better performance guarantees. More fundamentally, Proposition 3.3 leads naturally to the question of how much do the typical sets of different models overlap. The theory presented in this paper is neutral on this question which must be answered empirically and both possible answers have interesting implications. No overlap between typical sets means that each model can be associated with a unique statistical signature. On the other hand, independently trained models converging on a common set of typical strings, would be a notable linguistic phenomenon, though it is hard to speculate on what this might mean. The AEP is also just one example of the broader observation that generative AI models are stochastic machines subject to statistical laws and the discovery of other statistical regularities in the outputs of generative models is an important topic for future work.
Appendix A A Simple Markov Language Model
@@
Appendix B A Generalized Law of Large Numbers for Random Variables with Arbitrary Dependence
Consider a infinitely long string ${\mathbf{Y}}=[Y_{1}~{}Y_{2}~{}\dots~{}Y_{n}~{}\dots]$ consisting of a sequence of symbols $Y_{n}\in\mathcal{Y}$ generated stochastically from a finite dictionary $\mathcal{Y}$. We will denote the substring consisting of the first $N$ symbols as ${\mathbf{Y}}_{N}\equiv[Y_{1}~{}Y_{2}~{}\dots Y_{N}]$ and the joint distribution of the symbols $P({\mathbf{Y}}_{N})$. We make no assumptions about the distribution $P({\mathbf{Y}}_{N})$ from which the symbols $Y_{n}$ are generated. In particular, we do not assume that the symbols are independent or weakly dependent even asymptotically. We also do not assume stationarity i.e. the marginal distributions of symbols $Y_{n}$ and $Y_{m}$ can be completely different. We also do not assume any Markovian or conditional independence properties.
B.1 Conditional Means and Sample Means for Sequences of Dependent Random Variables
Consider a sequence of random variables ${\mathbf{Z}}=[Z_{1}~{}Z_{2}~{}\dots~{}Z_{n}~{}\dots]$ whose $n$’th element $Z_{n}$ is a function of the first $n$ symbols $Y_{1},Y_{2},\dots,Y_{n}$ i.e. $Z_{n}\doteq g_{n}({\mathbf{Y}}_{n})$ and $g_{n}:\mathcal{Y}^{n}\rightarrow\mathbb{R}$ are functions that we will assume to be uniformly bounded i.e. $|g_{n}({\mathbf{Y}}_{n})|\leq G<\infty,~{}\forall n\in\mathbb{Z}^{+},{\mathbf{Y}}_{n}\in\mathcal{Y}^{n}$.
We define the conditional mean and conditional variance of the random variables $Z_{n}$:
$\mu_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)g_{n}({\mathbf{Y}}_{n})\equiv\operatorname*{{E}}\left[Z_{n}|{\mathbf{Y}}_{n-1}\right]$ (26)
$V_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)\left(g_{n}({\mathbf{Y}}_{n})-\mu_{n}^{({\mathbf{Y}})}\right)^{2}\equiv\operatorname*{{E}}\left[Z_{n}^{2}|{\mathbf{Y}}_{n-1}\right]-\left(\mu_{n}^{({\mathbf{Y}})}\right)^{2}$ (27)
where the distribution $p^{({\mathbf{Y}})}_{n}(y)\doteq\Pr(Y_{n}=y|{\mathbf{Y}}_{n-1})$. Since the random variables $Z_{n}$ are discrete-valued and bounded, they have bounded moments for any distribution:
$|\mu_{n}^{({\mathbf{Y}})}|\leq G,~{}$ $V_{n}^{({\mathbf{Y}})}\leq G^{2}$ (28)
A more formal statement. Let $\mathcal{F}_{n}$ be the $\sigma$-algebra defined by ${\mathbf{Y}}_{n}\equiv[Y_{1}~{}Y_{2}~{}\dots~{}Y_{n}]$ i.e. $\mathcal{F}_{n}$ is a collection that includes all (mathematically reasonable) events defined on the first $n$ symbols. Then the random variables $Z_{n}$ as we have defined them are defined by the condition that $Z_{m}$ is $\mathcal{F}_{m}$-measurable, and $\mu_{m}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{m}\big{|}\mathcal{F}_{m-1}\right]$.
We need some intermediate results before presenting our main result. First, we define the partial sums of the sequence $Z_{n}$ and conditional means:
$S_{n}$ $\doteq\sum_{m=1}^{n}Z_{m},~{}\mu_{S_{n}}^{({\mathbf{Y}})}\doteq\sum_{m=1}^{n}\mu_{m}^{({\mathbf{Y}})}$ (29)
Clearly, $\operatorname*{{E}}\left[Z_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{n}^{({\mathbf{Y}})}\right]$ and $\operatorname*{{E}}\left[S_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{S_{n}}^{({\mathbf{Y}})}\right]$.
Lemma B.1.
The variance of the zero-mean random variable $S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}$ can be written as sums of contributions from each of the $Z_{m}$’s. Specifically, we have:
$V_{S_{n}}$ $\doteq\operatorname*{{E}}\left[\left(S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right)^{2}\right]=\sum_{m=1}^{n}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]$ (30)
Proof.
We will use the method of induction to prove (30). For $n=1$, (30) holds trivially. Now, let’s assume (30) holds for $n=l$ i.e. assume that:
$V_{S_{l}}$ $\equiv\sum_{m=1}^{l}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]$ (31)
Now consider $S_{l+1}\equiv S_{l}+Z_{l+1}$. Its variance can be written as:
$V_{S_{l+1}}$ $=\operatorname*{{E}}\left[\left(S_{l}+Z_{l+1}-\mu_{S_{l}}^{({\mathbf{Y}})}-\mu_{l+1}^{({\mathbf{Y}})}\right)^{2}\right]$ (32)
$=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]-2\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]$ (33)
We will now show that the last term in (33) is zero. We apply the law of iterated expectations to write:
$\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]=\operatorname*{{E}}_{{\mathbf{Y}}_{l}}\left[\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\big{|}{\mathbf{Y}}_{l}\right]\right]$ (34)
$=\operatorname*{{E}}_{{\mathbf{Y}}_{l}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\operatorname*{{E}}\left[\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\big{|}{\mathbf{Y}}_{l}\right]\right]\equiv 0$ (35)
where we used the fact that $S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}$ is a deterministic function of ${\mathbf{Y}}_{l}$ and $\mu^{({\mathbf{Y}})}_{l+1}\equiv\operatorname*{{E}}\left[Z_{l+1}|{\mathbf{Y}}_{l}\right]$.
Remark. An alternative way to arrive at (35) is to treat $\mu_{l+1}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{l+1}\big{|}{\mathbf{Y}}_{l}\right]$ as an estimate of $Z_{l+1}$ given ${\mathbf{Y}}_{l}$, and then use the famous Orthogonality Principle to argue that the estimation error $\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)$ must be uncorrelated with any function of ${\mathbf{Y}}_{l}$. More formally, $\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)$ is $\mathcal{F}_{l}$-measurable and is therefore orthogonal to $\left(Z_{l+1}-\operatorname*{{E}}\left[Z_{l+1}|\mathcal{F}_{l}\right]\right)$.
Using (35), we can now simplify (33) to:
$V_{S_{l+1}}$ $=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]\equiv\sum_{m=1}^{l+1}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]$ (36)
where we used the induction assumption (31). We have shown that if (30) holds for $n=l$, then (30) must also hold for $n=l+1$, thus completing the induction. ∎
Lemma B.2.
The sequence $\frac{1}{n}\left(S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right)$ converges to zero in probability:
$\lim_{n\rightarrow\infty}\Pr\left(\frac{1}{n}\left|S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~{}\forall\epsilon>0$ (37)
Proof.
Using Lemma B.1, we can write:
$V_{S_{n}}$ $=\sum_{m=1}^{n}\operatorname*{{E}}\left[V_{m}^{({\mathbf{Y}})}\right]\leq nG^{2}$ (38)
using the simple bounds in (28). The variance of $\frac{1}{n}\left(S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}\right)$ is simply $\frac{1}{n^{2}}V_{S_{n}}\leq\frac{G^{2}}{n}\rightarrow 0$ as $n\rightarrow\infty$. Equation (37) now follows immediately using the Chebyshev Inequality (Cohen, 2015). ∎
Lemma B.2 states that the sample means of the two random sequences $\{Z_{n}\}$ and $\{\mu^{({\mathbf{Y}})}_{n}\}$ must be close to each other asymptotically. Since we have made very few assumptions about the random variables involved, we cannot assume that either of the sequences $\{Z_{n}\}$ or $\{\mu^{({\mathbf{Y}})}_{n}\}$ themselves converge in any sense to an asymptotic limit. Neither can we assume that their sample means separately converge to a limit.
With an additional assumption stating the existence of one such limit, we can establish a Weak Law of Large Numbers generalized to its most extreme i.e. a statement that the sample mean of a sequence of random variables $\{Z_{n}\}$ with arbitrary dependence converges asymptotically to a suitably defined average quantity in probability.
Proposition B.1.
A Weak Law of Large Numbers. If the average of the conditional means converges to a limit i.e. $\mu^{({\mathbf{Y}})}\equiv\lim\limits_{n\rightarrow\infty}\frac{1}{n}\sum\limits_{m=1}^{n}\mu_{m}^{({\mathbf{Y}})}$ exists, then the sequence $\frac{1}{n}S_{n}$ converges in probability to $\mu^{({\mathbf{Y}})}$:
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=1}^{n}Z_{n}$ $\to_{p}\mu^{({\mathbf{Y}})}~{}~{}\mathrm{or}~{}\lim_{n\to\infty}\Pr\left(\left|\frac{1}{n}S_{n}-\mu^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~{}\forall\epsilon>0$ (39)
Proof.
Combining Lemma B.2 with the observation $\frac{1}{n}\mu_{S_{n}}^{({\mathbf{Y}})}\rightarrow\mu^{({\mathbf{Y}})}$ completes the proof. ∎
B.2 Discussion
Proposition B.1 is remarkable because unlike traditional versions of the LLN, it makes no assumptions whatsoever about the dependence or correlation between the symbols $Y_{n}$. It does not even require for instance that the correlation between $Y_{1}$ and $Y_{n}$ should decrease as $n$ becomes large.
The two extreme cases of strong dependence and complete independence for the symbols $Y_{n}$ turn out to be two simple special cases for our proposed LLN:
  • If the symbols $Y_{n}$ are independent, Proposition B.1 simply reduces to the classical weak LLN.
  • If the symbols are perfectly correlated, the $Z_{m}$ become deterministic functions of the first token, and the LLN then becomes a trivial observation.
Appendix C Supplemental Material
C.1 Additional Experimental Context and Results
In this work, all experiments were ran on a high-performance computing server running Linux (Ubuntu). AMD EPYC 7413 24-Core Processor, 48 logical cores, 96 GB RAM, four NVIDIA A100-SXM4-80GB GPUs. On this setup, each run of a text generation (i.e., running GPT-2 to generate Nmax = 512 tokens, Llama-8B to generate Nmax = 1024 tokens and Nmax = 10240 tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds to minutes, with each experiment taking about 90 minutes to run in the case of 10,240 steps. We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.3.
Figure 5 shows additional results similar to Fig. 2 where we naively varied the seed value of the random number generator from $209-214$ for Llama-8B. We present these to show that Fig. 2’s results are representative of the general case and were not cherry-picked. In all cases, $K=100$.
(a) Seed=209.
(b) Seed=210.
(c) Seed=211.
(d) Seed=212.
(e) Seed=213.
(f) Seed=214.
Figure 5: Log-perplexity and empirical entropy for Llama 3.1 of strings generated by the same model with top-100 sampling.